OpenStudy (anonymous):

which t-chart matches the equation y = x^2 - 2? a. x y -2 2 -1 -1 0 -2 1 -1 2 2 b. x y -2 -6 -1 -3 0 -2 1 -1 2 2 c. x y -2 2 -1 -1 0 -2 1 -3 2 6 d. x y -2 2 -1 1 0 -2 1 1 2 2

OpenStudy (anonymous):

@AravindG @amoodarya @Abhishek619 @awkwardpanda @Brad1996 @blackops2luvr @BlvckNight @bettylee @Callisto @Compassionate @CGGURUMANJUNATH @chucky00 @douglaswinslowcooper @Dobby1 @DullJackel09 @shelbygt520 @SolomonZelman @Spacelimbus @jagatuba @Jadeishere @Faris_Waleed @futureCollegeStudent @nannaberry @NadiaElhaj_1 @NotTheBrightest @Luigi0210 @LazyBoy @Jadeishere would any of you know :)

OpenStudy (dulljackel09):

Ima block you now

OpenStudy (anonymous):

okay :)

OpenStudy (jadeishere):

I think it's A but knowing me im probably wrong

OpenStudy (anonymous):

im with jade a

OpenStudy (amoodarya):

put x into it check y like this y = x^2 - 2 x=-2 -->y=4-2=2 x=-1 --->y+1-4=-3 can you go on ?

OpenStudy (anonymous):

I say A

OpenStudy (jadeishere):

yayI feelsmart

OpenStudy (anonymous):

its a

OpenStudy (amoodarya):

the aim is not to get answer ! but main goal is understanding